Calculate the amount of chlorine gas evolved during the electrolysis of molten potassium 

chloride on passing a current of 0.5 A for two hours through it.


1
Expert's answer
2021-03-29T05:18:23-0400

I= 0.5A, t=2hrs= 2x3600s=7200s

Q=It= 0.5x7200=3600C= 3600/96500= 0.037F

The half reaction occurring at the anode is

2Cl- ---> Cl2 + 2e-

2F liberates 1mol of Cl2

0.037F will liberate 1/2 x 0.037= 0.0186mole of Cl2